OpenStudy (anonymous):

Help please! Let z=-5sqrt3 / 2 + 5/2 i and w=1+sqrt3. a. Convert z and w to polar form b. Calculate zw using De Moivre’s Theorem c. Calculate (z / w) using De Moivre’s Theorem

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

\[z=-\frac{5\sqrt3}{2}+\frac{5}{2}i=5\left(-\frac{\sqrt3}{2}+\frac{1}{2}i\right)\] If you let \(z=a+bi\), then \(a=r\cos\theta\) and \(b=r\sin\theta\), so \[\begin{cases}r\cos\theta=5\left(-\dfrac{\sqrt3}{2}\right)\\r\sin\theta=5\left(\dfrac{1}{2}\right)\end{cases}~~\Rightarrow~~r=5,~\theta=\frac{5\pi}{6}\] So, \(z=5\left(\cos\dfrac{5\pi}{6}+i\sin\dfrac{5\pi}{6}\right)\). Is \(w=1+\sqrt3\) or is it \(1+\sqrt3 i\)?

OpenStudy (anonymous):

@SithsAndGiggles w=1+sqrt3i sorry and for the z it is -5 not negative the whole equation

OpenStudy (anonymous):

So is my \(z\) right or wrong? I'm not sure what you said.

OpenStudy (anonymous):

your z is wrong. It is just -5

OpenStudy (anonymous):

As in\[z=-5=-5(1+0i)~?\]

OpenStudy (anonymous):

\[z \frac{ -5\sqrt{3} }{ 2 } + \frac{ 5 }{ 2 }i\]

OpenStudy (anonymous):

Uh ... that's exactly what I had the first time.

OpenStudy (anonymous):

you had the - in front of the entire fraction. It is only supposed to be in front of the 5

OpenStudy (anonymous):

It's the same thing:\[-\frac{1}{2}=\frac{-1}{2}\]

OpenStudy (anonymous):

oh crap sill me. Ok so lets work on a.

OpenStudy (anonymous):

silly me

OpenStudy (anonymous):

Okay, well you have the trig form of \(z\). Let \(\large w=se^{i\phi}\), then \[\begin{cases}1=s\cos\phi\\ \sqrt3=s\sin\phi\end{cases}~~\Rightarrow~~s^2(\cos^2\phi+\sin^2\phi)=1^2+(\sqrt3)^2~~\Rightarrow~~s^2=4\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Rightarrow~~s=2,~\phi=\frac{\pi}{3}\] So, \(w=2\left(\cos\dfrac{\pi}{3}+i\sin\dfrac{\pi}{3}\right)\).

OpenStudy (anonymous):

Now, when you multiply two complex numbers in polar form, it's the same as multiplying the radii and adding the angles. In other words, \[\large z\cdot w=\left(re^{i\theta}\right)\left(se^{i\phi}\right)=rse^{i(\theta+\phi)}\] Division is a corresponding operation, where the radii are divided and the angles are subtracted: \[\large \frac{z}{w}=\frac{re^{i\theta}}{se^{i\phi}}=\frac{r}{s}e^{i(\theta-\phi)}\]

OpenStudy (anonymous):

so what are the polar forms @SithsAndGiggles would w be 2(cos pi/3 + isinpi/3) ?

OpenStudy (anonymous):

It's all there, yes.

OpenStudy (anonymous):

Ok, if those are the polar forms how exactly do I multiply and divide. I see what you said, but I'm not sure where to place everything

OpenStudy (anonymous):

If \(z=5\left(\cos\dfrac{5\pi}{6}+i\sin\dfrac{5\pi}{6}\right)\) and \(w=2\left(\cos\dfrac{\pi}{3}+i\sin\dfrac{\pi}{3}\right)\), then \[z\cdot w=(5\cdot2)\left(\cos\left(\frac{5\pi}{6}+\frac{\pi}{3}\right)+i\sin\left(\frac{5\pi}{6}+\frac{\pi}{3}\right)\right)\] Dividing is just as easy, just plug everything in.

OpenStudy (anonymous):

would wolfram calculate this properly? I don't have a calculator near me

OpenStudy (anonymous):

Are you telling me you don't know how to add fractions by hand? You probably don't need to find approximate values.

OpenStudy (anonymous):

is 10(-sqrt3/2 - i/2) the answer for the multiplication?

OpenStudy (anonymous):

Yes. Are you supposed to convert back from polar for the final answer?

OpenStudy (anonymous):

I'm not sure. There is nothing that says I must

OpenStudy (anonymous):

I know you said plug everything in for dividing, but which signs do I switch? All the + signs because / signs?

OpenStudy (anonymous):

\[\frac{z}{w}=\frac{5}{2}\left(\cos\left(\frac{5\pi}{6}-\frac{\pi}{3}\right)+i\sin\left(\frac{5\pi}{6}-\frac{\pi}{3}\right)\right)\]

OpenStudy (anonymous):

I got 5i/2 for this one

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

Ok, thanks a lot! I do have another question, so I'm not sure if you can help or not

OpenStudy (anonymous):

Ask away

OpenStudy (anonymous):

Here or new question?

OpenStudy (anonymous):

Go ahead and make it a new question, in case I can't help someone else may be able to.